OpenStudy (anonymous):

Choose the correct product of (6x + 2)2. 36x2 − 4 36x2 + 4 36x2 − 24x + 4 36x2 + 24x + 4

hartnn (hartnn):

\(\huge(a+b)^2=a^2+2ab+b^2 \\ \text{just put a=6x and b=2, what u get ?}\)

OpenStudy (anonymous):

how did you get that? These confuse me so much.. (6x+2)^2 = 6x^2 + 2(12) + b^2

hartnn (hartnn):

a= 6x a^2 = (6x)^2 = 6^2 x^2 = 36x^2 got this ?

OpenStudy (anonymous):

aaaah, its just confussing....

hartnn (hartnn):

which part ?

jimthompson5910 (jim_thompson5910):

Here's an alternate route: (6x + 2)^2 (6x + 2)(6x + 2) 6x*(6x + 2) + 2*(6x + 2) Now distribute. Then combine like terms.